OpenStudy (anonymous):

Mom pays $50.15 at a restaurant that included the meal and an 18% tip. What was the price of the meal alone? I know the answer is $42.50 but I don't know how to do the calculation.

OpenStudy (anonymous):

41.12 + 18% is $48.52. The Final amount paid for the bill is $50.15. Right?

OpenStudy (anonymous):

42.50 = 18% (7.65) = 50.15. I just don't know how to get the 42.50 number to start with.

OpenStudy (anonymous):

\[P(1.18)=50.15\] \[P=50.15\div 1.18\]

OpenStudy (anonymous):

Bingo! Thanks Satellite73!

OpenStudy (anonymous):

Cross multiply and keep everything in proportion. For example: x/50.18 = 18/100 (solve for x). x comes out to be roughly 9.03 which is the COST of the tip. Simply subtract that from your grand total. 50.15 - 9.03 = 41.12

OpenStudy (anonymous):

you CANNOT do it by finding 18% of the total and subtracting, because 18% of the total is bigger than 18% of the original bill. it is a division problem, not a subtraction

OpenStudy (anonymous):

i repeat, you CANNOT do it by finding 18% of 50.15

OpenStudy (anonymous):

What was the method you use satellite. I understand it completely, but why 1.18?

OpenStudy (anonymous):

to increase a number by 18% you multiply by 1.18 so in this case you have to divide by 1.18 because you don't know the original number, but rather you know the answer

OpenStudy (anonymous):

luis rivera wrote it too as long as you know that \[x+.18x=1.18x\]

OpenStudy (anonymous):

Thanks Satellite73 and Luis Rivera! For me, dividing by the 1.18 is the easiest method. Thanks to both for taking the time.